# Collision-Free Trajectory Optimization in Cluttered Environments with Sums-of-Squares Programming

### v1

@@ Abstract
In this work, we propose a trajectory optimization approach for robot navigation in cluttered 3D environments. We represent the robot’s geometry as a semialgebraic set defined by polynomial inequalities such that robots with general shapes can be suitably characterized. To address the robot navigation task in obstacle-dense environments, we exploit the free space directly to construct a sequence of free regions, and allocate each waypoint on the trajectory to a specific region. Then, we incorporate a uniform scaling factor for each free region, and formulate a Sums-of-Squares (SOS) optimization problem that renders the containment relationship between the robot and the free space computationally tractable. The SOS optimization problem is further reformulated to a semidefinite program (SDP), and the collision-free constraints are shown to be equivalent to limiting the scaling factor along the entire trajectory. In this context, the robot at a specific configuration is tailored to stay within the free region. Next, to solve the trajectory optimization problem with the proposed safety constraints (which are implicitly dependent on the robot configurations), we derive the analytical solution to the gradient of the minimum scaling factor with respect to the robot configuration. As a result, this seamlessly facilitates the use of gradient-based methods in efficient solving of the trajectory optimization problem. Through a series of simulations and real-world experiments, the proposed trajectory optimization approach is validated in various challenging scenarios, and the results demonstrate its effectiveness in generating collision-free trajectories in dense and intricate environments populated with obstacles.
@@ I Introduction
Over the past decade, substantial progress has been made in trajectory optimization for a diverse range of applications in robotics [1, 2, 3]. At its core, it aims to optimize for a collision-free trajectory that adheres to motion constraints while minimizing a task-specific cost function. Recent accomplishments in numerical optimization have facilitated effective solutions for nonlinear programs with various general constraints and objectives, resulting in improved performance across numerous robotic systems. Nonetheless, the inherent complexity and nonconvexity of the collision-free configuration space still pose significant challenges in effectively modeling safety constraints, particularly when confronted with intricate and densely populated obstacle layouts. Moreover, accurately formulating collision avoidance constraints often necessitates considering the geometry of the robot and its surrounding environment, which can be difficult to capture explicitly amidst complex robot and obstacle configurations. Consequently, there is a growing demand for advanced trajectory optimization techniques that can effectively address these challenges, substantially facilitating safe and efficient navigation of robots in complex and cluttered environments.
In this paper, we present a trajectory optimization approach that generates collision-free motions for robots with general geometries in intricate and cluttered environments. Specifically, it is achieved by enforcing the containment relationship between the robot and the collision-free workspace over the optimization horizon. We delineate the free space by decomposing it into a series of polytopic regions and constructing a graph to encode their connectivity information. Subsequently, we search for a sequence of free regions with the shortest edge cost that connects the start and goal configurations, based on which each waypoint is allocated to one specific free region. To render the safety condition tractable, we define a uniform scaling factor for the polytopic free region and formulate a Sums-of-Squares (SOS) optimization problem to solve for the minimal scaling factor that certifies the containment relationship between the robot and the scaled free region. It is worth noting that in the formulation of the safety constraints, the geometries of the robots are represented by sets of polynomial inequalities, thereby facilitating the application of this methodology to robots with general shapes. Our proposed SOS optimization problem is further recast as a semidefinite program (SDP). Moreover, the gradient of the scaling factor with respect to the robot configuration is subsequently derived and integrated with gradient-based trajectory optimization algorithm, resulting in an efficient pipeline that propels the trajectory towards the safe and optimal direction with swift and continuous convergence. Lastly, through extensive evaluations in a variety of challenging scenarios, our proposed framework demonstrates its feasibility and efficacy in generating safe and efficient trajectories, and the capability for flexible self-pose adjustment to adapt to complex and cluttered environments is attained.
@@ II Related Works
As an indispensable ingredient in trajectory optimization, various approaches have been explored in the literature to formulate collision avoidance constraints in cluttered environments. One category involves exploiting the distance information between robots and obstacles. In [4], distance margins between robot and obstacles are explicitly represented and enforced as state constraints. Analogously, distance functions are used to derive control constraints based on control barrier functions (CBFs), which are then synthesized into the design of safety-critical controllers [5]. However, the distance between two general shapes is inherently implicit, inevitably leading to oversimplified geometries of the robots and obstacles. In [6, 7], equivalent conditions for the implicit CBF constraints are derived by examining the perspective of the dual problem for the optimization problem that solves for the minimum distance between two polytopic sets, resulting in a trajectory optimization problem with nonlinear and nonconvex constraints. Another recent work [8] defines cone representations for several commonly encountered geometry primitives and constructs minimum scaling problems for every pair of collision primitives for collision checking. Gradient information is then extracted and combined with gradient-based solvers for collision-free trajectory generation in the scenario composed of these primitives. Yet, formulating distance constraints for each robot-obstacle pair leads to exponentially increased computational burden in cluttered environments. Additionally, these approaches require prior knowledge of obstacle geometries, which hinders their implementations in environments with intricate obstacle configurations.
Alternatively, a series of works enforce the robot to stay within the obstacle-free workspace [9, 10, 11, 12]. However, these methods typically assume point or circular robot models due to the intractability of the containment relationship between sets with general geometries. In [13], robots with specific convex geometries are treated as a point, and collision-free trajectory is generated inside the high-dimensional configuration free space, whose decomposition is obtained using the alternating optimization framework proposed in [14]. Although this method achieves high performance for motion planning in obstacle-dense environment, the process of generating such decomposition in high-dimensional configuration space is especially time-consuming, particularly when dealing with obstacles with complex configurations. Subsequently, a safety-critical reactive controller has been proposed that leverages the SOS techniques and moment constraints [15]. This approach filters the nominal twist control, ensuring that the robots with general geometries stay within the locally extracted polytopic free region in the workspace. However, the one-step controller tends to generate conservative and myopic strategies, utilizing only one surrounding free region extracted with current sensor input.
@@ IV-A Waypoints Allocation on Graph of Free Regions
As shown in Fig. 1(a), in an environment with random obstacles, we first generate a family of polytopic free regions at some sampling points using the decomposition algorithm in [9], such that $\mathcal{F}$ can be tightly approximated. Given $\mathcal{Q}$, we construct an undirected graph $\mathcal{G}:=(\mathcal{V},\mathcal{E})$, where each region $\mathcal{Q}_{i}$ is assigned to the vertex $v_{i}\in\mathcal{V}$, and the vertex pair ($v_{i},v_{j}$) is connected with an edge $e_{i,j}\in\mathcal{E}$ if they have an overlapping area $\mathcal{Q}_{i,j}$. Note that the overlapping area of two of these polytopic free regions, i.e., the intersection of two polytopes, is still a polytope whose representation can be easily obtained by stacking the linear constraints defining the two sets. Edge $e_{i,j}$ is defined as the line segments that connect $\mathcal{Q}_{i}$, $\mathcal{Q}_{i,j}$, and $\mathcal{Q}_{j}$ sequentially through their geometrical centers, and edge cost $\ell_{e_{i,j}}$ is the length of the straight lines.
Subsequently, as shown in Fig. 1(b), we search for a sequence of free regions $\mathbb{P}^{*}_{v}$ connecting $\boldsymbol{q}_{s}$ and $\boldsymbol{q}_{g}$ with the minimal path cost by applying Dijkstra algorithm on $\mathcal{G}$:
[…]
We further define two additional edges $e^{*}_{s,0}=(\boldsymbol{q}_{s},v^{*}_{0})$ and $e^{*}_{p,g}=(v^{*}_{p},\boldsymbol{q}_{g})$ as the line segments connecting the start position to $C(\mathcal{Q}^{*}_{0})$ and $C(\mathcal{Q}^{*}_{p})$ to the goal position, respectively. Then, the edge sequence concatenating each vertex in $\mathbb{P}^{*}_{v}$ can be extracted accordingly:
@@
To render the safety constraints in (III) computationally tractable, we adopt a heuristic method to enforce the robot to stay inside a designated free region at each time stamp:
@@
Specifically, for a given optimization horizon $T$, we first discretize the path in $\mathbb{P}^{*}_{e}$ by assuming a constant linear velocity to obtain $T+1$ discrete points, as demonstrated in Fig. 1(c). Then, each waypoint is assigned to the free region containing that point, and the points inside an overlapping area will always be assigned to the latter one according to the order in $\mathbb{P}^{*}_{e}$.
@@
Remark 1: The proposed pre-allocating method prevents introducing integer variables in the formulation of optimization problem (III), albeit at the expense of a degree of flexibility in trajectory assignments. However, given that the primary focus of this work is to ensure safety along the trajectory by exploring geometric considerations, and since mixed-integer programming is known to be NP-hard, the pre-allocating method proves to be a more pragmatic and efficient approach.
@@ IV-B Safety Constraints Formulation with SOS Programming
In this section, we introduce the transformation of the safety condition (2) into an SOS programming problem. Specifically, as illustrated in Fig. 2, we define $\mathcal{Q}_{\tau}(\alpha)$ as the uniformly scaled polytopic free region with a nonnegative scaling factor $\alpha$. Then, an optimization problem based on nonnegative polynomials and their SOS approximation is derived, which solves for the minimal scaling factor $\alpha$ such that the containment relationship between the robot and the scaled free region can be theoretically guaranteed. It is worthwhile to note that a solution always exists since the scaled free region can cover the whole Euclidean space $\mathbb{R}^{3}$ as $\alpha$ goes to infinity. In this sense, safety condition stated in (2) can be effectively enforced with $\alpha\leq 1$ along the trajectory without loss of generation.
@@ IV-B1 Set Representation with Polynomial Functions
We first introduce necessary basis for establishing the containment relationship between two semialgebraic sets defined with polynomial inequalities.
@@
For convenience, we will use $\boldsymbol{x}$ as the 3D coordinates in Euclidean space $\boldsymbol{x}=(x,y,z)$ in the remaining of this paper to describe the geometrical sets in $\mathbb{R}^{3}$ unless otherwise specified. The space occupied by robot $\mathcal{B}$ in its body frame $\{b\}$ is defined with $m$ polynomial inequalities:
@@
For a polytopic free region $\mathcal{Q}_{\tau}$ consisting of $r$ facets, we define it by stacking the linear constraints for each hyperplane into $F\in\mathbb{R}^{r\times 3}$ and $g\in\mathbb{R}^{r}$ in its own frame $q$:
${}^{q}\mathcal{Q}_{\tau}:=$$\{\boldsymbol{x}\in\mathbb{R}^{3}:g-F^{\top}\boldsymbol{x}\geq\boldsymbol{0}\}$
$=$ $\{\boldsymbol{x}\in\mathbb{R}^{3}:(g^{\top},-F^{\top})^{\top}[\boldsymbol{x}]_{1}\geq\boldsymbol{0}\},$
where $\boldsymbol{0}$ denotes the zero vector in $R^{r}$, and $[\boldsymbol{x}]_{1}=(1,x,y,z)$ is the monomial vector of degree 1 in $x$. The origin of $\{q\}$ is located at $C({}^{q}\mathcal{Q}_{\tau})$ with the axis aligning to the world frame $\{w\}$, as demonstrated in Fig. 2. In this sense, the uniformly scaled free region from $\{q\}$ can be represented as:
@@
which simply functions as a ratio on the constant term $g$.
@@ IV-B2 Certified Containment Relationship with Minimum Scaling
Denote $f^{\mathcal{Q}_{\tau}}$ the vector of the $r$ defining polynomials in ${}^{b}\mathcal{Q}_{\tau}(\alpha,\boldsymbol{q})$. The minimal scaling problem certifying the set containment relationship (4) is then re-stated in terms of the nonnegative polynomials cone on ${}^{b}\mathcal{W}_{\mathcal{B}}$:
@@
s.t. $f_{i}^{\mathcal{Q}_{\tau}}\in\mathcal{P}({}^{b}\mathcal{W}_{\mathcal{B}})$ (5)
@@
Here, $f^{\mathcal{Q}_{\tau}}_{i}$ represents the $i$th row of the polynomial vector $f^{\mathcal{Q}_{\tau}}$, which corresponds to the linear constraint defining the $i$th facet of the polytope. In the context of polynomial optimization, Putinar Positivestellensatz[17] provides us the theoretical guarantee to approximate $\mathcal{P}({}^{b}\mathcal{W}_{\mathcal{B}})$ with its quadratic module (Qmod) such that (IV-B2) can be described explicitly. The $2k$ truncation of Qmod for the defining polynomials $f$ in ${}^{b}\mathcal{W}_{\mathcal{B}}$ can be written as:
$\mbox{Qmod}[f]_{2k}:=\{\sigma_{0}+f_{1}\sigma_{1}+\ldots+f_{m}\sigma_{m}:\sigma_{0},\ldots,\sigma_{m}\in\Sigma[\boldsymbol{x}]\},$
with $\sigma_{0}\in\Sigma[\boldsymbol{x}]\cap\mathbb{R}[\boldsymbol{x}]_{2k}$ and $\sigma_{i}\in\Sigma[\boldsymbol{x}]\cap\mathbb{R}[\boldsymbol{x}]_{2\lfloor k-\deg(f_{i})/2\rfloor}$ for all $i=1,2,\dots,m$.
@@
s.t. $f^{\mathcal{Q}_{\tau}}_{i}=\sigma_{i,0}+\sum_{j=1}^{m}f_{j}\cdot\sigma_{i,j},$ (6a)
@@
s.t. ${\rm coef}(f^{\mathcal{Q}_{\tau}}_{i})_{l}-\sum_{j=0}^{m}\langle A_{(i,l),j},X_{i,j}\rangle=0,$ (7a)
@@
Remark 2: Since we have designated each waypoint on the trajectory $\mathcal{X}$ to a specific free region in Sec. IV-A, and considering that the set representation ${}^{b}\mathcal{W}_{\mathcal{B}}$ will not change with different robot configurations, it is possible to pre-compute all the matrix coefficients $A$, and therefore it greatly speeds up the online solving process.

### v2

# Collision-Free Trajectory Optimization in Cluttered Environments Using Sums-of-Squares Programming

@@ Abstract
In this work, we propose a trajectory optimization approach for robot navigation in cluttered 3D environments. We represent the robot’s geometry as a semialgebraic set defined by polynomial inequalities such that robots with general shapes can be suitably characterized. To address the robot navigation task in obstacle-dense environments, we exploit the free space directly to construct a sequence of free regions, and allocate each waypoint on the trajectory to a specific region. Then, we incorporate a uniform scaling factor for each free region, and formulate a Sums-of-Squares (SOS) optimization problem that renders the containment relationship between the robot and the free space computationally tractable. The SOS optimization problem is further reformulated to a semidefinite program (SDP), and the collision-free constraints are shown to be equivalent to limiting the scaling factor along the entire trajectory. In this context, the robot at a specific configuration is tailored to stay within the free region. Next, to solve the trajectory optimization problem with the proposed safety constraints, we derive a guiding direction for updating the robot configuration to decrease the minimum scaling factor by calculating the gradient of the Lagrangian at the primal-dual optimum of the SDP. As a result, this seamlessly facilitates the use of gradient-based methods in efficient solving of the trajectory optimization problem. Through a series of simulations and real-world experiments, the proposed trajectory optimization approach is validated in various challenging scenarios, and the results demonstrate its effectiveness in generating collision-free trajectories in dense and intricate environments populated with obstacles. Our code is available at: https://github.com/lyl00/minimum_scaling_free_region.
@@ I Introduction
Over the past decade, substantial progress has been made in trajectory optimization for a diverse range of applications in robotics[1, 2, 3]. At its core, it aims to optimize for a collision-free trajectory that adheres to motion constraints while minimizing a task-specific cost function. Recent accomplishments in numerical optimization have facilitated effective solutions for nonlinear programs with various general constraints and objectives. However, the inherent complexity and nonconvexity of the collision-free configuration space still pose significant challenges in effectively modeling safety constraints, particularly when confronted with intricate and densely populated obstacle layouts. Moreover, accurately formulating collision avoidance constraints often necessitates considering the geometry of the robot and its surrounding environment, which is difficult to capture explicitly amidst complex robot and obstacle configurations [4].
In this paper, we present a trajectory optimization approach that generates collision-free motions for robots with general geometries in intricate and cluttered environments. Specifically, it is achieved by enforcing the robots to stay within the collision-free workspace over the optimization horizon. We begin by approximating the free space by decomposing it into a series of polytopic regions and constructing a graph to capture how these regions interconnect. Subsequently, we search for a free region sequence and generate a path from the starting point to the destination that traverses through these free regions with the least cumulative distance. We formulate a Sums-of-Squares (SOS) optimization problem to solve for the minimum scaling factor that certifies the containment relationship between the robot and the scaled free region, considering its current configuration. The collision-free constraints are then enforced by limiting the scaling factor along the trajectory and the geometries of the robots are represented by sets of polynomial inequalities, thereby facilitating the application of this methodology to robots with general shapes in a unified framework. Moreover, the gradient of the scaling factor with respect to the robot configuration is subsequently approximated, which is effectively integrated with the augmented Lagrangian iterative linear quadratic regulator (AL-iLQR)[5], yielding a robust and efficient pipeline that steers the trajectory towards safety and optimality with rapid convergence. Lastly, through extensive evaluations in a variety of challenging scenarios, our proposed framework demonstrates its feasibility and efficacy in generating safe and efficient trajectories, and the capability for flexible self-pose adjustment adapting to complex and cluttered environments is attained.
@@ II Related Works
As an indispensable ingredient in trajectory optimization, various approaches have been explored in the literature to formulate collision avoidance constraints in cluttered environments. One category involves exploiting the distance information between robots and obstacles. In [6], distance margins between the robot and obstacles are explicitly represented and enforced as state constraints. However, the distance between two general shapes is inherently implicit, inevitably leading to oversimplified collision geometries and conservative maneuvers. In [7], collision avoidance between polytopic robots and obstacles is considered. It formulates a quadratic programming (QP) program that solves for the minimum distance between two polytopic sets. By examining the dual perspective of the QP, it derives explicit control barrier function (CBF) constraints and solves the nonlinear nonconvex optimization problem directly using IPOPT[8]. This method introduces numerous additional dual variables based on the optimization step size. Another recent work [9] defines cone representations for several common geometry primitives and constructs a minimum scaling problem for every pair of collision primitives for collision checking. However, this approach requires prior knowledge of obstacle geometries and restricts the robot’s shape to predefined primitives. Besides, different primitive pairs necessitate distinct optimization problems, which hinders their implementation in environments with complex obstacle configurations. Additionally, these distance-based methods rely on formulating distance constraints for each robot-obstacle pair, which significantly increases the computational burden in cluttered environments. Alternatively, some works have attempted to approximate obstacle-free workspaces with parametric free regions, such as polytopic free regions in Euclidean space [10] and elliptical safe corridors parameterized along a path [11]. In [10], the robot is constrained to stay within the polytopic free regions. However, this method assumes a point robot model due to the intractability of the containment relationship between sets with general geometries. In [12], the collision-free trajectory is generated in the high-dimensional configuration free space for robots with specific kinematic chains, whose convex decomposition is obtained using the alternating optimization framework proposed in [13]. While this approach represents an innovative effort to map high-dimensional, obstacle-free spaces, the process of achieving such decomposition proves to be notably time-intensive.
Numerous nonlinear programming (NLP) solvers have been developed to deal with trajectory optimization problems with nonlinear system dynamics[1, 8]. Among these, sequential convex programming (SCP) stands out as a robust strategy, which effectively approximates solutions through a series of convex subproblems. For instance, sequential quadratic programming (SQP) serves as the backbone algorithm in prominent tool such as TrajOpt[14]. Despite these advancements, the computational complexity has historically been a barrier to the broader application of SQP methods. In contrast, indirect methods like the iterative linear quadratic regulator (iLQR), offer advantages over these direct approaches, including reduced memory usage and enhanced computational speed. Extensions have been developed based on the barrier method[15], distributed optimization [16], and augmented Lagrangian techniques [17] to overcome the limitations of DDP-based methods in accommodating various constraints. Recently, ALTRO stands out as a significant advancement[5]. By incorporating iLQR within an augmented Lagrangian framework, ALTRO addresses both state and input constraints, demonstrating rapid convergence, numerical stability, and outperforming several benchmark solvers across a range of motion-planning challenges.
@@ IV-A Waypoints Allocation on Graph of Free Regions
As shown in Fig. 1(a), in an environment with random obstacles, we first generate a family of polytopic free regions by sampling points in $\mathcal{F}$ and using the decomposition algorithm in [10] till the portion of remaining free grids is below a threshold (1%-5%). Then, we construct an undirected graph $\mathcal{G}:=(\mathcal{V},\mathcal{E})$ from the generated free regions $\mathcal{Q}$. Specifically, each region $\mathcal{Q}_{i}$ represents a vertex (node) $v_{i}\in\mathcal{V}$. Two nodes are connected with an edge $e_{i,j}$ if regions $i$ and $j$ intersect sufficiently. Denote the overlapping area as $\mathcal{Q}_{i,j}$, and we represent each edge $e_{i,j}$ as the line segments that connect $\mathcal{Q}_{i}$, $\mathcal{Q}_{i,j}$, and $\mathcal{Q}_{j}$ sequentially through their geometrical centers, and edge cost $\ell_{e_{i,j}}$ is the length of the straight lines.
Subsequently, we search for a sequence of free regions $\mathbb{P}^{*}_{v}$ connecting $\boldsymbol{q}_{s}$ and $\boldsymbol{q}_{g}$ with the minimal cumulative edge length shown as the dark green regions in Fig. 1(b)) by applying Dijkstra algorithm on $\mathcal{G}$:
$\mathbb{P}^{*}_{v}=\left\{\{\boldsymbol{q}_{s},\mathcal{Q}^{*}_{0},\mathcal{Q}^{*}_{1},\ldots,\mathcal{Q}^{*}_{p},\boldsymbol{q}_{g}\}\left|\begin{array}[]{c}\mathcal{W}\left(\boldsymbol{q}_{s}\right)\subseteq\mathcal{Q}^{*}_{0},\\ \mathcal{W}\left(\boldsymbol{q}_{g}\right)\subseteq\mathcal{Q}^{*}_{p}\end{array}\right.\right\}.$
Two additional edges $e^{*}_{s,0}=(\boldsymbol{q}_{s},v^{*}_{0})$ and $e^{*}_{p,g}=(v^{*}_{p},\boldsymbol{q}_{g})$ as the line segments connecting the start position to $C(\mathcal{Q}^{*}_{0})$ and $C(\mathcal{Q}^{*}_{p})$ to the goal position, respectively. Then, the edge sequence concatenating each vertex in $\mathbb{P}^{*}_{v}$ can be extracted accordingly:
@@
For a given optimization horizon $T$, we discretize the path along the edge sequence $\mathbb{P}^{*}_{e}$ by assuming a constant linear velocity to obtain $T+1$ discrete waypoints. Then, we allocate each waypoint to the free region containing that point and the points inside an overlapping area will always be assigned to the latter one:
@@
Remark 1: We can determine the intersection by stacking the linear constraints defining the two polytopic sets. Consequently, the existence of overlap can be straightforwardly identified during the process of constructing the graph.
@@ IV-B Safety Constraints Formulation with SOS Programming
In this section, we introduce the transformation of the safety condition (2) into an SOS programming problem. Specifically, as illustrated in Fig. 2, we define $\mathcal{Q}_{\tau}(\alpha)$ as the uniformly scaled polytopic free region with a nonnegative scaling factor $\alpha$. Then, an optimization problem based on nonnegative polynomials cone is derived (Equation (IV-B2)) and solved to optimal via its SOS relaxation (Equations (6-7)). It solves for the minimal scaling factor $\alpha$ such that the containment relationship between the robot and the scaled free region can be certified. In this sense, safety can be guaranteed by enforcing the scaling factor $\alpha\leq 1$ along the trajectory without loss of generation. It is worthwhile to note that the minimum scaling factor $\alpha$ is unique and bounded since the scaled free region can cover the whole Euclidean space $\mathbb{R}^{3}$ as $\alpha$ goes to infinity.
@@ IV-B1 Set Representation with Polynomial Functions
We first introduce the necessary basis for establishing the containment relationship between two semialgebraic sets defined with polynomial inequalities.
@@
We will use $\boldsymbol{x}$ as the 3D coordinates in Euclidean space $\boldsymbol{x}=(x,y,z)$ in the remaining of this paper to describe the geometrical sets in $\mathbb{R}^{3}$ unless otherwise specified. The space occupied by robot $\mathcal{B}$ in its body frame $\{b\}$ is defined with $m$ polynomial inequalities:
@@
For a polytopic free region $\mathcal{Q}_{\tau}$ consisting of $r$ facets, we define it by stacking the linear constraints for each hyperplane into $F\in\mathbb{R}^{r\times 3}$ and $g\in\mathbb{R}^{r}$ in its own frame $q$. In this sense, the uniformly scaled free region from $\{q\}$ can be represented as:
@@
where $\alpha$ simply functions as a ratio on the constant term $g$.
@@ IV-B2 Certified Containment Relationship with Minimum Scaling
Then, the minimal scaling problem certifying the set containment relationship (4) at given $\boldsymbol{q}$ is then re-stated in terms of the nonnegative polynomials cone on ${}^{b}\mathcal{W}_{\mathcal{B}}$:
@@
s.t. $f_{i}^{\mathcal{Q}_{\tau}}(\alpha)\in\mathcal{P}({}^{b}\mathcal{W}_{\mathcal{B}})$ (5)
@@
Here, $f^{\mathcal{Q}_{\tau}}_{i}(\alpha)$ represents the $i$th defining linear constraints in ${}^{b}\mathcal{Q}_{\tau}(\alpha,\boldsymbol{q})$. Certifying the nonnegativity of a polynomial is generally intractable; therefore, we further introduce the quadratic module (Qmod) given by $f$:
$\mbox{Qmod}[f]:=\{\sigma_{0}+f_{1}\sigma_{1}+\ldots+f_{m}\sigma_{m}:\sigma_{0},\sigma_{1},\ldots,\sigma_{m}\in\Sigma[\boldsymbol{x}]\}.$
Theorem 1.
[19] Suppose ${}^{b}\mathcal{W}_{B}$ is Archimedean [^1] The set ${}^{b}\mathcal{W}_{B}$ is said to be Archimedean if there exists $\rho\in\mbox{Qmod}[f]$ such that the set given by $\rho(x)\geq 0$ is compact. Note that if ${}^{b}\mathcal{W}_{B}$ is compact, then one can always choose a sufficiently large number $N$ and add a redundant ball constraint $N-\sum_{i=1}^{n}x_{i}^{2}\geq 0$ to ${}^{b}\mathcal{W}_{B}$ such that the archimedean condition holds.. If a polynomial $p(\boldsymbol{x})>0$ for all $\boldsymbol{x}\in{{}^{b}\mathcal{W}_{B}}$, then $p\in\textup{Qmod}[f]$.
The above theorem is known as Putinar’s Positivstellensatz, which guarantees that all positive polynomials on $P(^{b}\mathcal{W}_{B})$ belong to $\mbox{Qmod}[f]$. Moreover, let the $2k$ truncation of Qmod, denoted as $\mbox{Qmod}[f]_{2k}$, be the subset of Qmod where the degree of each element is no greater than $2k$, i.e., $\sigma_{0}\in\Sigma[\boldsymbol{x}]\cap\mathbb{R}[\boldsymbol{x}]_{2k}$ and $\sigma_{i}\in\Sigma[\boldsymbol{x}]\cap\mathbb{R}[\boldsymbol{x}]_{2\lfloor k-\deg(f_{i})/2\rfloor}$ for all $i=1,2,\dots,m$. Then, for all $p\in P(^{b}\mathcal{W}_{B})$ and $\varepsilon>0$ ($\varepsilon$ can be 00 if $p>0$ on $P(^{b}\mathcal{W}_{B})$), there exists a degree $k$ such that $p+\varepsilon\in\mbox{Qmod}[f]_{2k}$. We refer to [20, 21] for more details on the convergence rate of Putinar’s Positivstellensatz.
@@
s.t. $f^{\mathcal{Q}_{\tau}}_{i}(\alpha)=\sigma_{i,0}+\sum_{j=1}^{m}f_{j}\cdot\sigma_{i,j},$ (6a)
@@
s.t. ${\rm coef}(f^{\mathcal{Q}_{\tau}}_{i}(\alpha))_{l}-\sum_{j=0}^{m}\langle A_{(i,l),j},X_{i,j}\rangle=0,$ (7a)
@@
Remark 2: Under mild conditions, optimal solutions of (IV-B2) and (7) are identical for all $k$ that is sufficiently large [22]. Indeed, they are equal for all $k$ if we further assume that all the polynomial functions involved are SOS-convex [18]. Therefore, we can determine the minimum scaling factor $\alpha$ in (IV-B2) by solving the SDP (7).